Suppose scores of a standardized test are normally distributed and have a known population standard deviation of 8 points and an unknown population mean. A random sample of 25 scores is taken and gives a sample mean of 93 points.

Find the margin of error for a confidence interval for the population mean with a 98% confidence level.

z0.10 z0.05 z0.025 z0.01 z0.005
1.282 1.645 1.960 2.326 2.576

You may use a calculator or the common z values above.

  • Round the final answer to two decimal places.
Answer #1

Solution: Given that mean = 93, population standard deviation = 8, n = 25

98% Confidence level for Z = 2.326 (by using above table)

margin of error = Z*σ/sqrt(n)
= 2.326*8/sqrt(25)
= 3.7216
= 3.72 (rounded)